OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 A)6 B)−6 C)All real numbers D)No solution

OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 -3x-9 =-3x-3-5 -3x-9 =-3x-8 -3x+3x = -8+9 0 = 1 d) no solution

OpenStudy (anonymous):

thank you soo much